Harris Fabrics computes its plantwide predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 39,000 direct labor-hours would be required for the period’s estimated level of production. The company also estimated $512,000 of fixed manufacturing overhead cost for the coming period and variable manufacturing overhead of $4.00 per direct labor-hour. Harris’s actual manufacturing overhead cost for the year was $743,714 and its actual total direct labor was 39,500 hours.

Compute the company’s plantwide predetermined overhead rate for the year. (Round your answer to 2 decimal places.)
Answer #1

Total  variable manufacturing overhead estimated=39,000*4=$156000

Total overhead estimated=Total  variable manufacturing overhead estimated+Total  fixed manufacturing overhead estimated

=156000+512,000

=$668000

Hence plantwide predetermined overhead rate=Total overhead estimated/Total estimated direct labor-hours

=$668000/39,000

=$17.13 per direct labor-hour(Approx).